OpenStudy (anonymous):

express number 8620 in 2421 code

OpenStudy (e.mccormick):

2421 code is sort of a binary code, but the but values are different from normal binary. \(\begin{array}{|c|c|c|c|c|}\hline \text{Spot}\rightarrow & 2 & 4 & 2 & 1\\ \downarrow \text{Decimal}\\ 0+0+0+0 = 0 & 0 & 0 & 0 & 0\\ \hline 0+0+0+1 =1 & 0 & 0 & 0 & 1\\ \hline 0+0+2+0 =2 & 0 & 0 & 1 & 0\\ \hline 0+0+2+1 =3 & 0 & 0 & 1 & 1\\ \hline 0+4+0+0 =4 & 0 & 1 & 0 & 0\\ \hline 2+0+1+1 =5 & 1 & 0 & 1 & 1\\ \hline 2+4+0+0 = 6 & 1 & 1 & 0 & 0\\ \hline 2+4+0+1 = 7 & 1 & 1 & 0 & 1\\ \hline 2+4+2+0 = 8 & 1 & 1 & 1 & 0\\ \hline 2+4+2+1 = 9 & 1 & 1 & 1 & 1\\ \hline \end{array}\) As you can see, it uses simple addition. Now, where binary does powers of 2, 2421 uses those specific numbers in that order. So for each digit, you need to put the correct encoding from on the right side of the chart.